Consider a market with 1,000 identical consumers with the following individual demand function p = 3 − qd and 1,000 identical sellers with the following individual supply function p = qs −1
. a) Obtain the market demand and supply functions and represent them graphically. b) Find the equilibrium price and quantity.Consider a market with 1,000 identical consumers with the following individual demand function p = 3 − qd and 1,000 identical sellers with the following individual supply function p = qs −1
. a) Obtain the market demand and supply functions and represent them graphically. b) Find the equilibrium price and quantity.
1
Expert's answer
2021-10-04T16:11:27-0400

"d=\frac{3-p}{q}"


"s=\frac{p+1}{q}"

d=s


"\frac{3-p}{q}" ="\frac{p+1}{q}"


"\frac{-2p+2}{q}=0"


p=1

q=2


d=1

s=1